OpenStudy (anonymous):

what is the ph of 1*10^-9 molar hcl

OpenStudy (anonymous):

ph = 9.0 :)

OpenStudy (anonymous):

it's an acid solution and pH in lower than 7.your answer is not correct. am i right?

OpenStudy (anonymous):

The trouble here is that 10^-9 molar HCl produces 10^-9 hydrogen ions, but that is significantly smaller than the 10^-7 hydrogen ions that are present from the water the HCl is dissolved in. So for an exact answer you have to consider the Ka of water and the expression for the total hydrogen ion concentration. But it's easy to estimate and say the pH is going to be just a little less than 7.

OpenStudy (aaronq):

Tech1 is right, add them up to see \([H^+]=1.0*10^{-9}+1.0*10^{-7}=1.01*10^{-7}~M\) \(pH=-log(1.01*10^{-7}~M)=6.9956\approx 7\)